Week of 230320

WLCG Operations Call details

  • The connection details for remote participation are provided on this agenda page.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to the wlcg-scod list (at cern.ch) to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • local: Julia (WLCG), Maarten (ALICE, WLCG), Panos (WLCG)
  • remote: Kate (DB, chair), David (IN2P3), Peter (ATLAS), Steve (computing), Xavier (KIT), Darren (RAL), Christoph (CMS), Alessandro (CNAF), Doug (BNL), Andrew (TRIUMF), Borja (monitoring), Dave (FNAL)

Experiments round table:

  • ATLAS reports (raw view) -
    • Issues
      • NDGF-T1 please check GGUS:160875
      • Alma9 Diffie-Hellman with keys less than 2048 bits in length are not accepted with the default policy. Two sites affected: INFN-GENOVA and INFN-LECCE. LEGACY crypto policy doesn't help.

  • CMS reports (raw view) -
    • While working with CERN EOS team on token access we discovered token utilities adding newline characters at the end of the token into token files;
      • xrootd doesn't handle this well
      • SAM tests changed to strip out control characters

    • ASGC Stratum-1 out of date since March 12 (GGUS:160909)
      • Due to very limited WAN bandwidth
      • Causing problems for 3 ALICE sites in Asia
    • Sites will finally be informed of the ALICE token configuration details this week

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • EGI:
  • FNAL: dCache intervention and storage downtime most probably tomorrow. Last tests ongoing.
  • IN2P3: NTR
  • JINR:
  • KISTI:
  • KIT: NTR
  • NL-T1:
  • NRC-KI:
  • OSG:
  • PIC: On Tuesday April 4th, PIC will be in complete scheduled downtime due to the building's yearly electrical maintenance. PIC will be in OUTAGE from 04-04-2023 04:00 [PIC local time] until 04-04-2023 23:59 [PIC local time]. HTCondor will be stopped at the beginning of the downtime.
  • RAL: NTR
  • TRIUMF: Will begin updating our 3 squid servers to frontier-squid-5.7-2.1. GOCDB: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=33711

  • CERN computing services:
    • CMS IAM CRIC Integration Release Wednesday 22nd at 08:30 OTG:0076430
  • CERN storage services:
  • CERN databases: NTR
  • Monitoring:
    • Distributed final SiteMon availability/reliability reports for February 2023
  • Middleware: NTR
  • Networks: NTR
  • Security:


Edit | Attach | Watch | Print version | History: r16 < r15 < r14 < r13 < r12 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r16 - 2023-03-20 - MaartenLitmaath
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback